OpenStudy (anonymous):

Determine the solution set of (x + 1)^2 = 25. {-8, 6} {-8, 4} {-6, 4}

OpenStudy (anonymous):

(x+1)^2= x^2+2x+1 plug in each set.

OpenStudy (anonymous):

am i just using the first number of the set or????

OpenStudy (anonymous):

@AEB047

OpenStudy (whpalmer4):

\[(x+1)^2=25\]That means that (x+1) represents a number which when multiplied with itself gives you 25. I'm sure you know that 5 is one of those numbers. But there is another number as well, can you tell me what it is?

OpenStudy (anonymous):

-5?

OpenStudy (whpalmer4):

Yes. So that means we have two equations to solve for \(x\): \[x+1 = 5\]\[x+1=-5\]The two values you get are the solution set.

OpenStudy (anonymous):

oh wow thanks dood

OpenStudy (whpalmer4):

You betcha. Test out your values in the original equation to make sure they are correct.

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

i didnt know it was that simple